# Aligning LLMs for Multilingual Consistency in Enterprise Applications

### v1

@@ 1 Introduction
The demand for multilingual natural language processing systems that perform reliably and consistently across diverse languages has grown rapidly in global industries such as customer support, content delivery networks, and information retrieval. Despite advances in LLMs, significant performance disparities persist between high-resource languages (e.g., English) and low-resource languages (e.g., French, Arabic), primarily due to the pretraining data, favoring high-resource corpora Xu et al. (2024). This gap undermines user experience, restricts the accessibility of AI technologies, and limits operational effectiveness for organizations serving multilingual populations worldwide.
@@ 5.2 RQ2: Generalization to Unseen Languages
We evaluate generalization on four unseen languages (Thai, Vietnamese, Hungarian, Romanian). ORPO with batch-alignment (ours) consistently outperform baselines in these languages (Table 1); for example, Mixtral-8x22B and Llama-3-70B achieve accuracy gains of up to 4.1%. This demonstrates that our batch alignment strategy enhances intrinsic cross-lingual generalization beyond the languages seen during fine-tuning. Table 7 in Appendix B expands the per language break-up and confidence-interval results.
@@ A.1 Extended Related Works
Cross-Lingual Transfer and Preference Alignment. Cross-lingual knowledge transfer remains essential for robust multilingual LLMs. Zhu et al. (2024) assessed LLM translation capabilities, noting that although models like GPT-4 improve high-resource language performance, they still lag behind specialized supervised systems in low-resource scenarios. Fine-tuning has shown promise over in-context learning, particularly for low-resource languages (Asai et al., 2024). Additionally, Shaham et al. (2024) proposed minimal multilingual instruction tuning, demonstrating that even limited multilingual data can significantly enhance cross-lingual generalization and instruction following on unseen languages.
@@
Several very recent works propose complementary strategies.  Zhu et al., 2023 and  Lin et al., 2024 explore cross-lingual instruction tuning, leveraging explicit translation tasks and mixed datasets to boost non-English and cross-lingual abilities, with a particular emphasis on data allocation and scaling. (Dang et al., 2024) and (She et al., 2024) advance preference-based optimization for multilingual LLMs, using reward models, synthetic or human preference data, and in some cases external translation models to align non-English outputs with English anchors.
@@
Ensuring consistent multilingual reasoning and generation within RAG systems is a challenging problem, critical to enterprise applications which serve the same information across different languages to a global user-base Pattnayak et al. (2025a); Meghwani et al. (2025). The NoMIRACL benchmark (Thakur et al., 2023) was introduced to rigorously evaluate RAG model’s ability to maintain factual consistency across languages by pairing queries with relevant and irrelevant documents in multiple languages.  (Zhang et al., 2024) focus on penalizing English responses to non-English prompts, addressing output language consistency without using multilingual data for fine-tuning, while effective, challenges remain due to differences in topic distributions and retrieval noise across languages. Zhang et al. (2023) further demonstrated challenges faced by multilingual LLMs with code-switched inputs, emphasizing the need for specialized training to handle linguistically complex scenarios.
Recent benchmarks like BUFFET (Asai et al., 2024), MMMLU OpenAI (2024), MGSM Cobbe et al. (2021); Shi et al. (2022), LAMBADA Multilingual EleutherAI (2024); Paperno et al. (2016), PAWS-X Yang et al. (2019) and xSTREET (Li et al., 2024a) reveal that LLMs still struggle to maintain consistent reasoning across languages, indicating that fundamental challenges remain; motivating the need for alignment-focused methods.
@@
Multilingual Fine-tuning Strategies. Fine-tuning multilingual models for specific tasks like document understanding Agarwal et al. (2025b); Yin et al. (2024); Patel et al. (2024); Panda et al. (2025a), video understanding Agarwal (2021); Agarwal et al. (2025a); Thomas et al. (2025), clinical trials Pattnayak et al. (2025c), and accessible applications Panda et al. (2025c, b) has evolved rapidly in recent years along with applications requiring both multi-modal and multilingual capabilities from models Pattnayak et al. (2024). A key advancement is the use of adapter layers Houlsby et al. (2019), which enable task-specific fine-tuning without modifying the entire model, resulting in reduced computational overhead. AdapterFusion Pfeiffer et al. (2020) further extends this by fusing knowledge from multiple task-specific adapters, allowing for better transfer across tasks and languages.
@@
Alignment Strategies for Multilingual Models. Beyond retrieval, aligning multilingual models internally has gained interest, which we further study in Appendix A.1. Recent methods explore multi-task and preference-based fine-tuning to reduce language biases and improve consistency (Patel et al., 2025; Rafailov et al., 2024; Hong et al., 2024b; She et al., 2024; Yang et al., 2024; Zeng et al., 2025). Our approach advances this by proposing a batch-wise alignment strategy using semantically equivalent multilingual data within each training batch, enabling direct alignment of generation across languages during fine-tuning without dependence on external translation or retrieval components.
@@ A.3.1 Metrics for Evaluation
While accuracy is well-suited to our binary QA formulation, we acknowledge the value of complementary metrics (such as semantic similarity or human evaluation) for more nuanced assessments of multilingual alignment.

### v2

@@ 1 Introduction
The demand for multilingual natural language processing systems that perform reliably and consistently across diverse languages has grown rapidly in global industries such as customer support, content delivery networks, and information retrieval Dua et al. (2025a); Agarwal et al. (2024). Despite advances in LLMs, significant performance disparities persist between high-resource languages (e.g., English) and low-resource languages (e.g., French, Arabic), primarily due to the pretraining data, favoring high-resource corpora Xu et al. (2024). This gap undermines user experience, restricts the accessibility of AI technologies, and limits operational effectiveness for organizations serving multilingual populations worldwide.
@@ 5.2 RQ2: Generalization to Unseen Languages
We evaluate generalization on four unseen languages (Thai, Vietnamese, Hungarian, Romanian). ORPO with batch-alignment (ours) consistently outperform baselines in these languages (Table 1); for example, Mixtral-8x22B and Llama-3-70B achieve accuracy gains of up to 4.1%. This demonstrates that our batch alignment strategy enhances intrinsic cross-lingual generalization beyond the languages seen during training. Table 7 in Appendix B expands the per language break-up and confidence-interval results.
@@ A.1 Extended Related Works
Cross-Lingual Transfer and Preference Alignment. Cross-lingual knowledge transfer remains essential for robust multilingual LLMs Son et al. (2025). Zhu et al. (2024) assessed LLM translation capabilities, noting that although models like GPT-4 improve high-resource language performance, they still lag behind specialized supervised systems in low-resource scenarios. Fine-tuning has shown promise over in-context learning, particularly for low-resource languages (Asai et al., 2024). Additionally, Shaham et al. (2024) proposed minimal multilingual instruction tuning, demonstrating that even limited multilingual data can significantly enhance cross-lingual generalization and instruction following on unseen languages.
@@
Several very recent works propose complementary strategies.  Zhu et al., 2023 and  Lin et al., 2024 explore cross-lingual instruction tuning, leveraging explicit translation tasks and mixed datasets to boost non-English and cross-lingual abilities, with a particular emphasis on data allocation, extraction and scaling Panda et al. (2025b). (Dang et al., 2024) and (She et al., 2024) advance preference-based optimization for multilingual LLMs, using reward models, synthetic or human preference data Patel et al. (2024a); Panda et al. (2025c); Dua et al. (2025b), and in some cases external translation models to align non-English outputs with English anchors.
@@
Ensuring consistent multilingual reasoning and generation within RAG systems is a challenging problem, critical to enterprise applications which serve the same information across different languages to a global user-base Wasi et al. (2025); Pattnayak et al. (2025a); Meghwani et al. (2025); Singh (2023, 2021). The NoMIRACL benchmark (Thakur et al., 2023) was introduced to rigorously evaluate RAG model’s ability to maintain factual consistency across languages by pairing queries with relevant and irrelevant documents in multiple languages.  (Zhang et al., 2024) focus on penalizing English responses to non-English prompts, addressing output language consistency without using multilingual data for fine-tuning, while effective, challenges remain due to differences in topic distributions and retrieval noise across languages and data types Dua et al. (2025a); Singh (2022). Zhang et al. (2023) further demonstrated challenges faced by multilingual LLMs with code-switched inputs, emphasizing the need for specialized training to handle linguistically complex scenarios.
Recent benchmarks like BUFFET (Asai et al., 2024), MMMLU OpenAI (2024), MGSM Cobbe et al. (2021); Shi et al. (2022), BenchHub Kim et al. (2025), SEA-VL Cahyawijaya et al. (2025), LAMBADA Multilingual EleutherAI (2024); Paperno et al. (2016), PAWS-X Yang et al. (2019) and xSTREET (Li et al., 2024a) reveal that LLMs still struggle to maintain consistent reasoning across languages, indicating that fundamental challenges remain; motivating the need for alignment-focused methods.
@@
Multilingual Fine-tuning Strategies. Fine-tuning multilingual models for specific tasks like document understanding Agarwal et al. (2025b); Yin et al. (2024); Patel et al. (2024b); Panda et al. (2025a), video understanding Agarwal (2021); Agarwal et al. (2025a); Thomas et al. (2025), clinical trials Pattnayak et al. (2025c), and accessible applications Panda et al. (2025d, e, f) has evolved rapidly in recent years along with applications requiring both multi-modal and multilingual capabilities from models Pattnayak et al. (2024). A key advancement is the use of adapter layers Houlsby et al. (2019), which enable task-specific fine-tuning without modifying the entire model, resulting in reduced computational overhead. AdapterFusion Pfeiffer et al. (2020) further extends this by fusing knowledge from multiple task-specific adapters, allowing for better transfer across tasks and languages.
@@
Alignment Strategies for Multilingual Models. Beyond retrieval, aligning multilingual models internally has gained interest, which we further study in Appendix A.1. Recent methods explore multi-task and preference-based fine-tuning to reduce language biases and improve consistency (Patel et al., 2025a; Rafailov et al., 2024; Hong et al., 2024b; She et al., 2024; Yang et al., 2024; Zeng et al., 2025). Our approach advances this by proposing a batch-wise alignment strategy using semantically equivalent multilingual data within each training batch, enabling direct alignment of generation across languages during fine-tuning without dependence on external translation or retrieval components.
@@ A.3.1 Metrics for Evaluation
While accuracy is well-suited to our binary QA formulation, we acknowledge the value of complementary metrics (such as semantic similarity or human evaluation) for more nuanced assessments of multilingual alignment across diverse textual and visual tasks Agarwal et al. (2025c); Patel et al. (2025b).